# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 4.2 Nice 4-Tuples in $G(m,p,3)$
Let $[r_{2},r_{3},r_{4},r_{5}]$ be a nice 4-tuple in $G(m,1,3)$. Let $r_{5}$ be the unique reflection of Type 2 with nontrivial eigenvalue $\zeta^{-1}$ of order $m$, so we can write $\zeta^{-1}=\zeta_{m}^{-b}$ with $(-b,m)=1$. Let $d=order(\lambda)$. We know $d\mid m$ since $\lambda I$ is in the center $ZG(m,1,3)$. Taking determinants we get the equation
@@
Let $m$ be such that $\gcd(m,3)=1$, so we know $m-3$ is coprime to $m$. Taking $b=m-3$, we can consider the tuple $[s(1,2;0),s(2,3;1),s(2,3;-1),s(3;m-3)]$. Taking $X=\{s(1,2;b_{2}),s(2,3;b_{3}),s(3;-b)\}$ we see by Theorem 4.0.4 that $\langle X\rangle=G(m,1,3)$ since $\gcd(m-3,m)=1$. Taking the product, we get $r_{2}r_{3}r_{4}r_{5}=\lambda r_{1}^{-1}$ for some Type 1 reflection $r_{1}=s(1,2;-1)$ with $\lambda=\zeta_{m}$. ∎
@@
Let $[r_{2},r_{3},r_{4},r_{5}]$ be a nice 4-tuple in $G(m,p,3)$. Letting $r_{5}$ denote the unique reflection of Type 2, we know it must have order $q=p^{-1}m$ so we can denote its nontrivial eigenvalue as $\zeta=\zeta_{q}^{b}=\zeta_{m}^{pb}$ with $\gcd(b,m)=1$. Further, we have the product $r_{2}r_{3}r_{4}r_{5}=\lambda r_{1}^{-1}$ with $r_{1}\in G(m,m,3)\leq G(m,p,3)$, hence $\lambda I$ is in the center $ZG(m,p,3)$. Letting $d=order(\lambda)$ we see that since $|ZG(m,p,3)|=q\gcd(p,3)$, we must have that $d\mid q\gcd(p,3)$. Taking determinants, we get $\lambda^{3}=\zeta^{-1}$ and hence $q\mid d$ and $d\mid 3q$. If $d\nmid q$ then since $d\neq 3$ (becasue $\zeta^{-1}\neq 1$), and $d\mid q\gcd(p,3)$ we must have that $\gcd(p,3)=3$. This gives us two cases:
@@
In either case we get $\gcd(pb,\delta(X),m)=k$ so we conclude $k=1$. This also gives us that $\lambda$ must be of order $m$. Thus, we conclude nice 4-tuples can only exist in $G(m,3,3)$. In this case, we can write $\lambda=\zeta_{m}^{a}$ with $(a,m)=1$ and we still have $\zeta=\zeta_{m}^{3b}$ with $\gcd(b,m)=1$ for $q=p^{-1}m$. We can take our 4-tuple to be $[s(1,2;0),s(2,3;1),s(2,3;-1),s(3;m-3)]$, noting that since $3\mid m$ we also have that $\gcd(m-3,m)=3$. Here $\delta(X)=2$ so $\gcd(m-3,2,m)=1$ and we conclude this tuple is indeed nice. ∎
@@ 4.3 Nice 4-Tuples in $G(m,p,4)$
Suppose we had $r_{1}r_{2}r_{3}r_{4}=A$ such that $r_{i}$ generate $G$ and $A$ has eigenvalue $\lambda\neq 1$ of multiplicity 2. Since $\{r_{i}\}$ generate $G$, we know exactly three of them are of Type 1 and exactly one is of Type 2 with nontrivial eigenvalue an $m$-th root of unity, say $\zeta_{m}^{\prime}$. Note that since the three Type 1 reflections should correspond to transpositions that generate all of $S_{4}$, they must be three distinct transpositions with each of $\{1,2,3,4\}$ showing up at least once. This means their product cannot be a transposition. Using $\varphi:G\longrightarrow S_{4}$ we get that $\varphi(A)=\sigma$ for some 4-cycle $\sigma$ since the permutation must be odd and cannot be a transposition. Writing $A=[a_{1},a_{2},a_{3},a_{4}\mid\sigma]$ we can compute that the character polynomial of $A$ $p_{A}(x)$ must be
$p_{A}(x)=x^{4}-\zeta_{m}^{a_{1}+a_{2}+a_{3}+a_{4}}$
@@
The number of transitive ordered factorizations of $(12)(34)$ into four transpositions in $S_{4}$ is 96 [16]. Using a computer, we find that there are 12 factorizations up to simultaneous conjugation. We can reduce these 12 even further by eliminating factorizations that are related by the action of one of the generators of $B_{4}$, since this means they lie in the same braid group orbit. These preliminary computations give us three equivalence classes of factorizations, (which may still lie in the same braid group orbit) which we write with representatives:
@@
11 11 $=\det(A)=\lambda^{2}\lambda_{1}\lambda_{2}$
@@
$11=|\lambda|=\frac{|\lambda_{1}+\lambda_{2}|}{2}=\frac{|\lambda_{1}|+|\lambda_{2}|}{2}=$ 1
By properties of the triangle inequality, we get that $\lambda_{1}$ and $\lambda_{2}$ are colinear, hence $\lambda_{1}=\pm\lambda_{2}$. However, if $\lambda_{1}=-\lambda_{2}$ then the trace equation means $\lambda=0$ which is not possible, hence $\lambda_{1}=\lambda_{2}$. Now the trace equation tells us that $02(\lambda+\lambda_{1})=$ 0, hence $\lambda_{1}=-\lambda$. Next, we can use the determinant equation to see that $1=\lambda^{4}$ and so $\lambda$ must be a 4th root of unity. This gives us that the eigenvalues of $A$ are as described in case 1).
Further, when $\varphi(A)$ is a double transposition, say $A=[a_{1},a_{2},a_{3},a_{4}\mid(12)(34)]$, the characteristic polynomial is $p_{A}(x)=(x^{2}-\zeta_{m}^{a_{1}+a_{2}})(x^{2}-\zeta_{m}^{a_{3}+a_{4}})$. However, since we know the roots of these polynomials agree, we see that $\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{a_{3}+a_{4}}$, but we already knew that $\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{-a_{3}-a_{4}}$ since $m\mid a_{1}+a_{2}+a_{3}+a_{4}$. Thus we conclude that $\zeta_{m}^{a_{1}+a_{2}}=\pm 1$ and so $\zeta_{m}^{a_{1}}=\pm\zeta_{m}^{-a_{2}}$ and similarly, $\zeta_{m}^{a_{3}}=\pm\zeta_{m}^{-a_{4}}$ where the signs must agree for both pairs.
@@
Computing the characteristic polynomial we get $p_{A}(x)=(x-\zeta_{m}^{a_{4}})(x^{3}-\zeta_{m}^{a_{1}+a_{2}+a_{3}})$ and since the cube roots of $\zeta_{m}^{a_{1}+a_{2}+a_{3}}$ are all distinct, we know that for $\lambda$ to have multiplicity two, it must be one of the cube roots and $\lambda=\zeta_{m}^{a_{4}}$. Since $\zeta_{m}^{a_{1}+a_{2}+a_{3}}=\zeta_{m}^{-a_{4}}=\lambda^{-1}$, we see that $\lambda,\lambda_{1}$ and $\lambda_{2}$ must be the three distinct cube roots of $\lambda^{-1}$. Now we have that $\lambda^{3}=\lambda^{-1}$ and hence $\lambda^{4}=1$, proving that the eigenvalues of $A$ are as described in case 2). ∎
@@
If $A$ corresponds to a double transposition, then our nice 4-tuple is equivalent to one of the factorizations in Remark 4.3.2 and we may write $A=[a_{1},a_{2},a_{3},a_{4}\mid(12)(34)]$. From the first factorization and the proof of Lemma 4.3.4, we see that ±$\pm 1=\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{b_{3}-b_{4}}=\lambda^{2}$. In this case $\delta(X)=|b_{3}-b_{4}|$ which we know must be coprime to $m$ by Theorem 4.0.4, we get that $\lambda^{2}$ is a primitive $m$-th root so $m=2$. The same argument will also give that $m$ must be 2 if our nice tuple is equivalent to the second factorization in Remark 4.3.2. For the third possible factorization, we see that the cycle has four edges and so $\delta(X)=|b_{1}-b_{3}+b_{2}-b_{4}$ since the cycle corresponds to the permutation $(1324)$. But again we have that ±$\pm 1=\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{b_{1}+-b_{3}+b_{2}-b_{4}}$ so $\lambda^{2}$ is a primitive $m$-th root so $m=2$. If $m=4$, then we can take $r_{1}=s(1,4;-1),r_{2}=s(1,4;0),r_{3}=s(1,2;-1),r_{4}=s(2,3;0)$. Then $\delta(X)=1$ so these reflections generate $G(4,4,4)$. Further, we have
@@ 4.4 Nice 5-Tuples in $G(m,p,4)$
Searching for 5-tuples $G(m,p,4)$ so that the middle convolution will be rank 2 means we want the product to have an eigenvalue of multiplicity 3. A $4\times 4$ matrix with an eigenvalue of multiplicity 3 can be viewed as a scalar times a reflection, so we want to explore the existence of $[r_{1},r_{2},r_{3},r_{4},r_{5}]$ such that
@@
Note first that $G(m,m,4)\leq G(m,p,4)\leq G(m,1,4)$ so $r_{i}\in G(m,1,4)$ for $i<6$ and so $\lambda r_{6}^{-1}\in G(m,1,4)$. Since this matrix has an eigenvalue of multiplicity 3, we know $d=order(\lambda)$ is such that $d$ divides 3 of the degrees in $\deg G(m,1,4)=\{m,2m,3m,4m\}$. This tells us that $d\mid m$ as if it did not then $d\mid 3m-2m=m$. Since $\lambda$ has order dividing $m$, $\lambda I$ is in the center of $G(m,1,4)$ and thus we can conclude $r_{6}\in G(m,1,4)$, though not yet necessarily in $G(m,p,4)$. This changes our problem into studying 6 reflections in $G(m,1,4)$ satisfying
@@
The second and fourth equation give us $-b_{1}+b_{2}-b_{4}\equiv 2k2^{s-j}$ which is not coprime to $m$. Thus, this option for $\delta(X)$ does not work so one of the first three options are coprime to $m$ and hence $\lambda$ is of order $m$. Since $\lambda$ is a 4-th root of unity, we know $m=2$ or $m=4$.
A similar argument will give the same conclusion for the second factorization. For $m=4$, consider the tuple $[s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;1)]$. Then $\delta(X)=1$ and so these reflections will generate $G(4,4,4)$. The product of this tuple will be $i\cdot s(2,4;2)$ as this corresponds to the first factorization. For $m=2$, we can take the tuple $[s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;0)]$. Then $\delta(X)=1$ and so these reflections will generate $G(2,2,4)$. The product of this tuple will be $-1\cdot s(2,4;1)$ as this corresponds to the first factorization. ∎
@@ 4.5 Nice $n$-Tuples in $G(m,p,n)$ for $n\geq 5$
Fact: A connected graph of $n$ vertices and $n-1$ edges must be a tree.
@@
The product of $n-1$ transpositions is an $n$-cycle if and only if the graph associated to a set of transposition in $S_{n}$ is a tree, where the edge $(a,b)$ is only present if $(ab)$ is one of the $n-1$ transpositions.
@@
By induction: if the product of $n-1$ transpositions is the product of a disjoint $k$-cycle and an $(n-k)$-cycle, then each of these correspond to disjoint trees with $k-1$ and $(n-k-1)$ edges respectively, so the remaining edge must connect the two trees, resulting in a tree with $n-1$ edges, hence the product is an $n$-cycle. ∎
@@
In $G(m,1,n)$ any nice $n$-tuple consists of $n-1$ reflections of Type 1 and one reflection of Type 2, say $r_{n}$. Suppose we had $r_{1}r_{2}\cdots r_{n}=A$ with $[r_{1},\dots,r_{n}]$ a nice tuple. Then by the claim(s) above, the graph associated to the transpositions $\{\varphi(r_{i})\}$, with $\varphi(r_{n})=e$ must be a tree since the graph is connected by transitivity and therefore $\varphi(A)$ is an $n$-cycle. Thus the characteristic polynomial
@@
for some $m$-th root of unity $\zeta$ and therefore has $n$ distinct roots so no $A$ has no eigenvalues of multiplicity $n-2$ for $n>3$. ∎
@@
Let $\sigma=(x_{1}x_{2},\cdot x_{n})$ be an $n$-cycle in $S_{n}$ and $\tau=(y_{1}y_{2})$ be some transposition in $S_{n}$ with $y_{1}<y_{2}$. Letting $c$ denote the distance between $y_{1}$ and $y_{2}$ in $\sigma$ modulo the length of $\sigma$, we have that the product $\sigma\cdot\tau$ is the product of a disjoint $c$-cycle and a disjoint $(n-c)$-cycle.
@@
Note that since $\sigma$ is written in disjoint cycle notation the ‘distance’ between, i.e. number of digits between $y_{1}$ and $y_{2}$ read in the order of multiplication (e.g. left to right if multiplication is left to right) is well defined.
@@
If $y_{1}$ and $y_{2}$ are adjacent then $c=1$, so we are left with just an $n-1$ cycle which has exactly 1 fixed point when $n-1>1$ and 2 fixed points when $n=2$. ∎
@@
In $G(m,m,n)$, any nice tuple $[r_{1},r_{2},\dots,r_{n}]$ consists of $n$ reflections of Type 1. Again by transitivity, we know the graph associated to $\{\varphi(r_{i})\}$ is connected and since we only need $n-1$ edges to make a connected graph on $n$ vertices, we know there is some transposition we can remove so that the associated graph is still connected. By cyclic permutation of the $r_{i}$ via simultaneous conjugation, we can assume the redundant transposition corresponds to $r_{n}$, so $\varphi(A\cdot r_{n}^{-1})$ is an $n$-cycle by the claim above. This tells us that $\varphi(A)$ is the product of an $n$ cycle and a transposition. By the proofs above, we know $\varphi(A)$ is either an $(n-1)$-cycle or the product of a $c$-cycle and an $(n-c)$-cycle. So either the characteristic polynomial is
@@
We see that in both cases an eigenvalue of $A$ can have atmost multiplicity 2, so when $n>4$, there are no eigenvalues of multiplicity $n-2$, hence no such nice tuples can exist in $G(m,m,n)$. ∎

### v3

@@ 4.2 Nice 4-Tuples in $G(m,p,3)$
Let $r_{2},r_{3},r_{4},r_{5}]$ be a nice 4-tuple in $G(m,1,3)$. Let $r_{5}$ be the unique reflection of Type 2 with nontrivial eigenvalue $\zeta^{-1}$ of order $m$, so we can write $\zeta^{-1}=\zeta_{m}^{-b}$ with (b,m)=1(-b,m)=1. Let $d=order(\lambda)$. We know $d\mid m$ since $\lambda I$ is in the center $ZG(m,1,3)$. Taking determinants we get the equation
@@
Let $m$ be such that $\gcd(m,3)=1$, so we know m3m-3 is coprime to $m$. Taking b=m3b=m-3, we can consider the tuple [s(1,2;0),$s(2,3;1),$s(3;m3)][s(1,2;0),s(2,3;1),s(2,3;-1),s(3;m-3)]. Taking $X=\{s(1,2;b_{2}),s(2,3;b_{3}),s(3;-b)\}$ we see by Theorem 4.0.4 that $\langle X\rangle=G(m,1,3)$ since $\gcd(m-3,m)=1$. Taking the product, we get $r_{2}r_{3}r_{4}r_{5}=\lambda r_{1}^{-1}$ for some Type 1 reflection $r_{1}=s(1,2;-1)$ with $\lambda=\zeta_{m}$. ∎
@@
Let $r_{2},r_{3},r_{4},r_{5}]$ be a nice 4-tuple in $G(m,p,3)$. Letting $r_{5}$ denote the unique reflection of Type 2, we know it must have order $q=p^{-1}m$ so we can denote its nontrivial eigenvalue as $\zeta=\zeta_{q}^{b}=\zeta_{m}^{pb}$ with $\gcd(b,m)=1$. Further, we have the product $r_{2}r_{3}r_{4}r_{5}=\lambda r_{1}^{-1}$ with $r_{1}\in G(m,m,3)\leq G(m,p,3)$, hence $\lambda I$ is in the center $ZG(m,p,3)$. Letting $d=order(\lambda)$ we see that since $|ZG(m,p,3)|=q\gcd(p,3)$, we must have that $d\mid q\gcd(p,3)$. Taking determinants, we get $\lambda^{3}=\zeta^{-1}$ and hence $q\mid d$ and $d\mid 3q$. If $d\nmid q$ then since $d\neq 3$ (becasue $\zeta^{-1}\neq 1$), and $d\mid q\gcd(p,3)$ we must have that $\gcd(p,3)=3$. This gives us two cases:
@@
In either case we get $\gcd(pb,\delta(X),m)=k$ so we conclude $k=1$. This also gives us that $\lambda$ must be of order $m$. Thus, we conclude nice 4-tuples can only exist in $G(m,3,3)$. In this case, we can write $\lambda=\zeta_{m}^{a}$ with $(a,m)=1$ and we still have $\zeta=\zeta_{m}^{3b}$ with $\gcd(b,m)=1$ for $q=p^{-1}m$. We can take our 4-tuple to be [s(1,2;0),$s(2,3;1),$s(3;m3)][s(1,2;0),s(2,3;1),s(2,3;-1),s(3;m-3)], noting that since $3\mid m$ we also have that $\gcd(m-3,m)=3$. Here $\delta(X)=2$ so $\gcd(m-3,2,m)=1$ and we conclude this tuple is indeed nice. ∎
@@ 4.3 Nice 4-Tuples in $G(m,p,4)$
Suppose we had $r_{1}r_{2}r_{3}r_{4}=A$ such that $r_{i}$ generate $G$ and $A$ has eigenvalue $\lambda\neq 1$ of multiplicity 2. Since $\{r_{i}\}$ generate $G$, by Theorem 4.0.1 we know exactly three of them are of Type 1 and exactly one is of Type 2 with nontrivial eigenvalue an $m$-th root of unity, say $\zeta_{m}^{\prime}$. Note that since the three Type 1 reflections should correspond to transpositions that generate all of $S_{4}$, they must be three distinct transpositions with each of $\{1,2,3,4\}$ showing up at least once. This means their product cannot be a transposition. Using $\varphi:G\longrightarrow S_{4}$ we get that $\varphi(A)=\sigma$ for some 4-cycle $\sigma$ since the permutation must be odd and cannot be a transposition. Writing $A=[a_{1},a_{2},a_{3},a_{4}\mid\sigma]$ we can compute that the characteristic polynomial of $A$ must be
$char_{A}(x)=x^{4}-\zeta_{m}^{a_{1}+a_{2}+a_{3}+a_{4}}$
@@
The number of transitive ordered factorizations of $(12)(34)$ into four transpositions in $S_{4}$ is 96 [transfac]. Using a computer, we find that there are 12 factorizations up to simultaneous conjugation. We can reduce these 12 even further by eliminating factorizations that are related by the action of one of the generators of $B_{4}$, since this means they lie in the same braid group orbit. These preliminary computations give us three equivalence classes of factorizations, (which may still lie in the same braid group orbit) which we write with representatives:
@@
$1$ $=\det(A)=\lambda^{2}\lambda_{1}\lambda_{2}$
@@
$1=|\lambda|=\frac{|\lambda_{1}+\lambda_{2}|}{2}=\frac{|\lambda_{1}|+|\lambda_{2}|}{2}=1$
By properties of the triangle inequality, we get that $\lambda_{1}$ and $\lambda_{2}$ are colinear, hence $\lambda_{1}=\pm\lambda_{2}$. However, if $\lambda_{1}=-\lambda_{2}$ then the trace equation means $\lambda=0$ which is not possible, hence $\lambda_{1}=\lambda_{2}$. Now the trace equation tells us that $2(\lambda+\lambda_{1})=0$, hence $\lambda_{1}=-\lambda$. Next, we can use the determinant equation to see that $1=\lambda^{4}$ and so $\lambda$ must be a 4th root of unity. This gives us that the eigenvalues of $A$ are as described in case 1).
Further, when $\varphi(A)$ is a double transposition, say $A=[a_{1},a_{2},a_{3},a_{4}\mid(12)(34)]$, the characteristic polynomial is $char_{A}(x)=(x^{2}-\zeta_{m}^{a_{1}+a_{2}})(x^{2}-\zeta_{m}^{a_{3}+a_{4}})$. However, since we know the roots of these polynomials agree, we see that $\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{a_{3}+a_{4}}$, but we already knew that $\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{-a_{3}-a_{4}}$ since $m\mid a_{1}+a_{2}+a_{3}+a_{4}$. Thus we conclude that $\zeta_{m}^{a_{1}+a_{2}}=\pm 1$ and so $\zeta_{m}^{a_{1}}=\pm\zeta_{m}^{-a_{2}}$ and similarly, $\zeta_{m}^{a_{3}}=\pm\zeta_{m}^{-a_{4}}$ where the signs must agree for both pairs.
@@
Computing the characteristic polynomial we get $char_{A}(x)=(x-\zeta_{m}^{a_{4}})(x^{3}-\zeta_{m}^{a_{1}+a_{2}+a_{3}})$ and since the cube roots of $\zeta_{m}^{a_{1}+a_{2}+a_{3}}$ are all distinct, we know that for $\lambda$ to have multiplicity two, it must be one of the cube roots and $\lambda=\zeta_{m}^{a_{4}}$. Since $\zeta_{m}^{a_{1}+a_{2}+a_{3}}=\zeta_{m}^{-a_{4}}=\lambda^{-1}$, we see that $\lambda,\lambda_{1}$ and $\lambda_{2}$ must be the three distinct cube roots of $\lambda^{-1}$. Now we have that $\lambda^{3}=\lambda^{-1}$ and hence $\lambda^{4}=1$, proving that the eigenvalues of $A$ are as described in case 2). ∎
@@
If $A$ corresponds to a double transposition, then our nice 4-tuple is equivalent to one of the factorizations in Remark 4.3.2 and we may write $A=[a_{1},a_{2},a_{3},a_{4}\mid(12)(34)]$. From the first factorization and the proof of Lemma 4.3.4, we see that $\pm 1=\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{b_{3}-b_{4}}=\lambda^{2}$. In this case $\delta(X)=|b_{3}-b_{4}|$ which we know must be coprime to $m$ by Theorem 4.0.4, we get that $\lambda^{2}$ is a primitive $m$-th root so $m=2$. The same argument will also give that $m$ must be 2 if our nice tuple is equivalent to the second factorization in Remark 4.3.2. For the third possible factorization, we see that the cycle has four edges and so $\delta(X)=|b_{1}-b_{3}+b_{2}-b_{4}$ since the cycle corresponds to the permutation $(1324)$. But again we have that $\pm 1=\zeta_{m}^{a_{1}+a_{2}}=\zeta_{m}^{b_{1}+-b_{3}+b_{2}-b_{4}}$ so $\lambda^{2}$ is a primitive $m$-th root so $m=2$. If $m=4$, then we can take $r_{1}=s(1,4;-1),r_{2}=s(1,4;0),r_{3}=s(1,2;-1),r_{4}=s(2,3;0)$. Then $\delta(X)=1$ so these reflections generate $G(4,4,4)$. Further, we have
@@ 4.4 Nice 5-Tuples in $G(m,p,4)$
Searching for 5-tuples $G(m,p,4)$ so that the middle convolution will be rank 2 means we want the product to have an eigenvalue of multiplicity 3. A $4\times 4$ matrix with an eigenvalue of multiplicity 3 can be viewed as a scalar times a reflection, so we want to explore the existence of $r_{1},r_{2},r_{3},r_{4},r_{5}]$ such that
@@
Note first that $G(m,m,4)\leq G(m,p,4)\leq G(m,1,4)$ so $r_{i}\in G(m,1,4)$ for $i<6$ and so $\lambda r_{6}^{-1}\in G(m,1,4)$. Since this matrix has an eigenvalue of multiplicity 3, we know $d=order(\lambda)$ is such that $d$ divides 3 of the degrees in $\deg G(m,1,4)=\{m,2m,3m,4m\}$. This tells us that $d\mid m$ as if it did not then $d\mid 3m-2m=m$, which would be a contradiction. Since $\lambda$ has order dividing $m$, $\lambda I$ is in the center of $G(m,1,4)$ and thus we can conclude $r_{6}\in G(m,1,4)$, though not yet necessarily in $G(m,p,4)$. This changes our problem into studying 6 reflections in $G(m,1,4)$ satisfying
@@
The second and fourth equation give us $b_{1}+b_{2}-b_{4}\equiv 2k2^{s-j}$ which is not coprime to $m$. Thus, this option for $\delta(X)$ does not work so one of the first three options are coprime to $m$ and hence $\lambda$ is of order $m$. Since $\lambda$ is a 4-th root of unity, we know $m=2$ or $m=4$.
A similar argument will give the same conclusion for the second factorization. For $m=4$, consider the tuple [s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;1)][s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;1)]. Then $\delta(X)=1$ and so these reflections will generate $G(4,4,4)$. The product of this tuple will be $i\cdot s(2,4;2)$ as this corresponds to the first factorization. For $m=2$, we can take the tuple [s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;0)][s(1,2;1),s(1,3;0),s(1,2;0),s(2,3;1),s(2,4;0)]. Then $\delta(X)=1$ and so these reflections will generate $G(2,2,4)$. The product of this tuple will be $1\cdot s(2,4;1)$ as this corresponds to the first factorization. ∎
@@ 4.5 Nice $n$-Tuples in $G(m,p,n)$ for $n\geq 5$
Fact: A connected graph of $n$ vertices and n1n-1 edges must be a tree.
@@
The product of n1n-1 transpositions is an $n$-cycle if and only if the graph associated to the set of n1n-1 transpositions in $S_{n}$ is a tree, where the edge $(a,b)$ is present if and only if $(ab)$ is one of the n1n-1 transpositions.
@@
We first check the“only if” direction: if the product of n1n-1 transpositions is an $n$-cycle, the associated graph must be connected; since it has $n$ vertices and n1n-1 edges, the fact above implies that it is a tree. It remains to prove the“if” direction.
We proceed by induction. Assume the result holds for $k<n$. Given a tree with $n$ vertices, there must exist some vertex $b$ of degree one. Removing this vertex $b$ and its only edge $(a_{m},b)$ is still a tree, so by the induction hypothesis, we have that the product of the remaining n2n-2 transpositions is an (n1)(n-1)-cycle $(a_{1}\cdots a_{m}\cdots a_{n-1})$ with $a_{i}\neq b$ for all $i$. This tells us that the product $(a_{1}\cdots a_{m}\cdots a_{n-1})(a_{m}b)=(a_{1}\cdots a_{m}ba_{m+1}\cdots a_{n-1})$, which is an $n$-cycle. ∎
@@
In $G(m,1,n)$ any nice $n$-tuple consists of n1n-1 reflections of Type 1 and one reflection of Type 2, say $r_{n}$. Suppose we had $r_{1}r_{2}\cdots r_{n}=A$ with $r_{1},\dots,r_{n}]$ a nice tuple. Then by the claim(s) above, the graph associated to the transpositions $\{\varphi(r_{i})\}$, with $\varphi(r_{n})=e$ must be a tree since the graph is connected by transitivity and therefore $\varphi(A)$ is an $n$-cycle. Thus the characteristic polynomial
@@
for some $m$-th root of unity $\zeta$ and therefore has $n$ distinct roots so no $A$ has no eigenvalues of multiplicity n2n-2 for $n>3$. ∎
@@
Let $\sigma=(x_{1}x_{2}\cdots x_{n})$ be an $n$-cycle in $S_{n}$ and $\tau=(y_{1}y_{2})$ be some transposition in $S_{n}$ with $y_{1}<y_{2}$. Letting $c$ denote the distance between $y_{1}$ and $y_{2}$ in $\sigma$ modulo the length of $\sigma$, we have that the product $\sigma\tau$ is the product of a $c$-cycle and an (nc)(n-c)-cycle, which are disjoint.
@@
Note that since $\sigma$ is written in disjoint cycle notation the ‘distance’ between, i.e. number of elements between $y_{1}$ and $y_{2}$ read in the order of multiplication (e.g. left to right if multiplication is left to right) is well-defined.
@@
If $y_{1}$ and $y_{2}$ are adjacent then $c=1$, so we are left with just an n1n-1 cycle which has exactly 1 fixed point when n1>1n-1>1 and 2 fixed points when $n=2$. ∎
@@
In $G(m,m,n)$, any nice tuple $r_{1},r_{2},\dots,r_{n}]$ consists of $n$ reflections of Type 1. Again by transitivity, we know the graph associated to $\{\varphi(r_{i})\}$ is connected and since we only need n1n-1 edges to make a connected graph on $n$ vertices, we know there is some transposition we can remove so that the associated graph is still connected. By cyclic permutation of the $r_{i}$ via simultaneous conjugation, we can assume the redundant transposition corresponds to $r_{n}$, so $\varphi(A\cdot r_{n}^{-1})$ is an $n$-cycle by the claim above. This tells us that $\varphi(A)$ is the product of an $n$ cycle and a transposition. By the proofs above, we know $\varphi(A)$ is either an (n1)(n-1)-cycle or the product of a $c$-cycle and an (nc)(n-c)-cycle. So either the characteristic polynomial is
@@
We see that in both cases an eigenvalue of $A$ can have at most multiplicity 2, so when $n>4$, there are no eigenvalues of multiplicity n2n-2, hence no such nice tuples can exist in $G(m,m,n)$. ∎